# Universal quantum control over non-Hermitian continuous-variable systems

### v1

@@ Abstract
Although the control of non-Hermitian quantum systems has a growing interest for their nonunitary feature in the time evolution, the existing discussions are not more than two or three dimensions and heavily influenced by the singularity of the energy spectrum. We here develop a general theory to control an arbitrary number of bosonic modes governed by the time-dependent non-Hermitian Hamiltonian. It takes advantage of the gauge potential in the instantaneous frame rather than the energy spectrum of Hamiltonian. In particular, the dynamics of a general non-Hermitian continuous-variable system is analyzed in the instantaneous frame associated with time-dependent ancillary operators that are superpositions of the laboratory-frame operators and irrelevant to the original Hamiltonian. The gauge potential is determined by the unitary transformation between the time-dependent and stationary ancillary frames. The upper triangularization condition for the Hamiltonian’s coefficient matrix in the stationary ancillary frame enables two of the time-dependent ancillary operators to be nonadiabatic Heisenberg passages of the non-Hermitian system. The probability conservation of the system wavefunction can be restored at the end of these passages without artificial normalization. Our theory is exemplified with the perfect and nonreciprocal state transfers in a cavity magnonic system. The former holds for arbitrary initial states and is irrelevant to the parity-time symmetry of the Hamiltonian and the exceptional point of the spectra; and the latter is consistent with the unidirectional perfect absorbtion. Our work essentially extends the universal quantum control (UQC) theory to the non-Hermitian continuous-variable systems, providing a promising approach for their coherent control.
@@ I Introduction
Conventional control protocols about non-Hermitian systems are spectrum-dependent and typically under the Hamiltonian of parity-time ($\mathcal{PT}$) symmetry Bender and Boettcher (1998); Bender and Hook (2024). Compared to Hermiticity, $\mathcal{PT}$-symmetry of Hamiltonian is a wider criterion for determining whether the system’s eigen-spectra is real or complex. When the Hamiltonian commutes with the parity-time-inversion operator, i.e., $H,\mathcal{PT}]=0$, the system is featured with real eigenenergies. In the broken phase of $\mathcal{PT}$-symmetry, i.e., $H,\mathcal{PT}]\neq 0$, the system can have complex eigenenergies Bender and Boettcher (1998); Bender and Hook (2024). Recently, much effort has been devoted to the presence or absence of exceptional points (EPs) Heiss (2012). In both unbroken and broken phases of $\mathcal{PT}$ symmetry Ashida et al. (2020), EPs might emerge when some eigenvalues and eigenstates of system coalesce and the biorthogonal condition breaks down Brody (2013). The non-Hermitian Hamiltonian arises generally either from the energy or material exchange with the environment Metelmann and Clerk (2015); Harder et al. (2018); Grigoryan et al. (2018); Wang et al. (2019a) or from post-selection over no-quantum-jump trajectories Han et al. (2024). The current work aims to establish a universal theoretical framework for non-Hermitian continuous-variable systems, which is irrelevant to the energy spectrum and is not constrained by the increasing complexity with system size. It is a nontrivial extension of the universal quantum control (UQC) over closed systems Jin and Jing (2025a, b, c, d, e) that are isolated from the external environment.
@@
For the ideal control over closed bosonic systems, the Hermitian Hamiltonian preserves the state norm and the real-valued energy spectrum. However, the quantum systems are inevitably open in practical scenarios, such as photonic systems Guo et al. (2009); Regensburger et al. (2012); Ge and Türeci (2013); Peng et al. (2014); Hodaei et al. (2017); Zhang et al. (2020); Bergman et al. (2021), acoustic systems Zhu et al. (2014); Fleury et al. (2015); Christensen et al. (2016); Liu et al. (2018); Wang et al. (2019b); Tang et al. (2020), and cavity magnonic systems Zhang et al. (2017); Wang et al. (2019a); Zhang et al. (2019); Zhao et al. (2020); Qian et al. (2024); Yang et al. (2024). Their interaction with the environment leads to a breakdown of unitary evolution and a violation of probability conservation of the system wavefunction Ashida et al. (2020). Due to the cooperative coupling to a common reservoir, the non-Hermitian Hamiltonian is used to account for the gain or loss effects associated with individual bosonic modes Zhang et al. (2017, 2025) and the dissipative coupling between the system modes Metelmann and Clerk (2015); Harder et al. (2018); Grigoryan et al. (2018); Wang et al. (2019a). In particular, the loss effect on a cavity mode can be effectively converted into the gain effect by applying external microwave fields to its ports Zhang et al. (2017, 2025).
In the unbroken phase of $\mathcal{PT}$ symmetry and in the absence of EPs, perfect state transfer can be performed in the two- and three-level systems Ibáñez et al. (2011); Torosov et al. (2013, 2014), employing the time-dependent gain or loss rates of populations and the renormalization of the quantum trajectories. By tuning the Rabi frequency and detuning of the driving fields to dynamically encircle an EP, one can create Bell states Khandelwal et al. (2024) and realize the population transfer between two eigenstates Liu et al. (2021) in the nitrogen-vacancy-center system. Within the scattering-matrix framework, a directional quantum amplifier for two photonic modes Metelmann and Clerk (2015) can be achieved when their coherent coupling is precisely matched by their engineered dissipative coupling. Nevertheless, the stringent experimental requirements Ibáñez et al. (2011); Torosov et al. (2013, 2014) and artificial normalization Daley et al. (2009); Uzdin et al. (2012); Ibáñez and Muga (2014); Ashida and Ueda (2018); Dóra and Moca (2020) will become inaccessible when extended to larger-scale systems. The universal quantum control framework Jin and Jing (2025a, b, f, c, g, d, e), however, exploits the gauge potential supported by the differential manifold of rotating frames, on which an energy-spectrum-independent theory can be proposed for the control over a general non-Hermitian bosonic systems.
@@
In this paper, the Heisenberg equation about the non-Hermitian and time-dependent bosonic systems is prescribed by the instantaneous ancillary operators. Transformation from the time-dependent to the time-independent (stationary) ancillary frames introduces a gauge potential, by which the triangularization condition of the Hamiltonian’s coefficient matrix enables the exact solutions of the Heisenberg equation for two of the ancillary operators. In contrast to the conventional treatments of non-Hermitian systems Daley et al. (2009); Uzdin et al. (2012); Ibáñez and Muga (2014); Ashida and Ueda (2018); Dóra and Moca (2020), our theory permits a deterministic evolution toward a desired target state, inherently conserving probability at the end of the passage without artificial normalization. In a cavity magnonic system with gain or loss effects on individual modes and a dissipative coupling between them, it is found that arbitrary initial state in the cavity mode can be perfectly transferred to the magnon mode and perfect absorbtion can occur if the initial excitation is prepared in the magnon mode. In other words, the cavity-magnonic system under our UQC theory can be used as a promising unidirectional perfect absorber Ramezani et al. (2014); Longhi (2015); Jin et al. (2016); Ramezani et al. (2016); Huang et al. (2017); Jin and Song (2018); Xu and Jin (2024) at the operational level.
@@
The rest of this paper is structured as follows. In Sec. II, we introduce a general theory for solving the Heisenberg equation of the bosonic operators under an arbitrary non-Hermitian and time-dependent Hamiltonian. The Hamiltonian’s coefficient matrix can be triangularized in a stationary frame and then it suffices to activate two ancillary operators to become rapid adiabatic passages. In Sec. III.1, our theory is used to analyze a non-Hermitian cavity-magnonic system; and in Sec. III.2, we construct the relevant Heisenberg passages that in general do not conserve the state-norm. Section IV.1 demonstrates that arbitrary state prepared in the cavity mode can be perfectly transferred to the magnon mode irrespective to the parity-time-symmetry of Hamiltonian and the presence or absence of exceptional points. In contrast, Sec. IV.2 shows that a unidirectional perfect absorbtion occurs when the interested state is initially prepared in the magnon mode. The entire work is summarized in Sec. V. Appendix A provides a brief recipe for constructing the ancillary operators as well as the unitary transformation between the time-dependent and stationary ancillary frames. Appendix B details the derivation of the non-Hermitian Hamiltonian based on the Lindblad master equation for the cavity-magnonic system.
@@ II General framework
where $\vec{a}\equiv(a_{1},a_{2},\cdots,a_{N})$ and $\vec{a}^{\dagger}=(a_{1}^{\dagger},a_{2}^{\dagger},\cdots,a_{N}^{\dagger})$ are row operator vectors, and $H^{a}(t)$ is an $N\times N$ time-dependent coefficient matrix with Hermitian-conjugate $H^{a}(t)]^{\dagger}$. $H^{a}(t)\neq[H^{a}(t)]^{\dagger}$. The superscript $T$ denotes the matrix transposition, also transforming a row vector to its corresponding column vector. In both discrete- and continuous-variable systems, solving the time-dependent Schrödinger equation with Hermitian or non-Hermitian Hamiltonian remains a challenging task. Here, we extend our UQC theory Jin and Jing (2025a, b, f, c, g, d, e) to the non-Hermitian bosonic systems, providing a fundamental framework for partially solving Eqs. (1a) and (1b).
Without loss of generality, we start with the system dynamics in the ket space, as given by Eq. (1a) with $H(t)$ in Eq. (2). In the spirit of universal quantum control for Hermitian bosonic systems Jin and Jing (2025e), the system dynamics can be described in the ancillary representation associated with a completed set of time-dependent ancillary operators $\mu_{k}(t)$’s, $1\leq k\leq N$. $\{\mu_{k}(t)\}$ are connected with the laboratory-frame bosonic operators $\{a_{k}\}$ by an $N\times N$ unitary transformation matrix $\mathcal{M}^{\dagger}(t)$:
@@
A specific unitary transformation matrix $\mathcal{M}^{\dagger}(t)$ and the relevant formation of ancillary operators are provided in appendix A. Due to Eq. (3) and the unitary feature of $\mathcal{M}^{\dagger}(t)$, one can find that the ancillary operators $\mu_{k}(t)$’s satisfy the canonical communication relation, i.e., $\mu_{j}^{\dagger},\mu_{k}(t)]=-\delta_{jk}$. Using Eq. (3), the time-dependent Hamiltonian (2) can be formulated as
@@
is the Hamiltonian’s coefficient matrix in the representation of time-dependent ancillary bases $\mu_{k}(t)|{\rm vac}\rangle$, where $|{\rm vac}\rangle$ represents the vacuum state.
To proceed, we consider the rotation from the time-dependent ancillary operators to the time-independent or stationary ancillary operators, i.e., $\vec{\mu}_{t}\rightarrow\vec{\mu}_{0}$ with $\vec{\mu}_{0}=[\mu_{1}(0),\mu_{2}(0),\cdots,\mu_{N}(0)]$. Such a rotation can be achieved by the unitary transformation $\mathcal{V}(t)$ as $\mathcal{V}^{\dagger}(t)\mu_{k}(t)\mathcal{V}(t)\rightarrow\mu_{k}(0)$, where $\mathcal{V}(t)$ can be determined by $\mathcal{M}^{\dagger}(t)$ in Eq. (3) (see appendix A for details). Generally, in the rotating frame with respect to $\mathcal{V}(t)$, we have
@@
Consequently, the time-dependent Schrödinger equations in Eq. (1) are transformed as
@@
Still it remains difficult to determine the system dynamics by solving the evolution operators in Eq. (9), owing to the noncommunity of $H_{\rm rot}(t)$ at distinct moments. And it is not appropriate to apply the commutation condition about the Hamiltonian’s coefficient matrix and the projection operator in the representation of time-independent ancillary modes established for the Hermitian bosonic network Jin and Jing (2025e) to Eq. (6) due to the fact that the non-Hermitian coefficient matrix $H^{\mu}(t)-\mathcal{A}(t)$ is generally non-diagonalizable, as $H^{\mu}_{km}(t)\neq H^{\mu}_{mk}(t)$ for $k\neq m$.
Main result.— we prove that the upper triangularization of the coefficient matrix of $H_{\rm rot}(t)$ is a sufficient condition for activating two useful Heisenberg-picture passages, which is only relevant to the differential manifold rather than the spectral characteristics. Specifically, the upper triangularization condition for the coefficient matrix $H^{\mu}(t)-\mathcal{A}(t)$ is given by Jin and Jing (2025g); Ju et al. (2024)
@@
where $T$ means matrix transposition and $\Pi^{k}$ with $1\leq k\leq N$ is the projection operator or an $N\times N$ matrix defined by $\Pi^{k}_{jm}=\delta_{jk}\delta_{mk}$.
@@
The dynamics of an arbitrary operator $\tilde{\mathcal{O}}_{S}$ can be obtained by $\mathcal{O}_{H}(t)=V_{\rm rot}^{\dagger}(t)\tilde{\mathcal{O}}_{S}U_{\rm rot}(t)$ according to the non-Hermitian Heisenberg equation Miao and Xu (2016) with the evolution operators given by Eq. (9). Specifically under the Hamiltonian (11), the dynamics of the ancillary operators $\mu_{1}^{\dagger}(0)$ and $\mu_{N}(0)$ are found to be decoupled from the others and therefore can be derived as $v_{1}^{\dagger}(t)\equiv V_{\rm rot}^{\dagger}(t)\mu_{1}^{\dagger}(0)U_{\rm rot}(t)$ and $v_{N}(t)\equiv V_{\rm rot}^{\dagger}(t)\mu_{N}(0)U_{\rm rot}(t)$. In particular, we have
@@
Equations (16) and (17) indicate that the ancillary operators $\mu_{1}^{\dagger}(t)$ and $\mu_{N}(t)$ are activated as useful passages toward the desired target state. For the systems initially prepared in the states $F[\mu_{1}^{\dagger}(0)]|{\rm vac}\rangle$ and $F[\mu_{N}(0)]|{\rm vac}\rangle$, they will evolve to the states $F[e^{if_{1}(t)}\mu_{1}^{\dagger}(t)]|{\rm vac}\rangle$ and $F[e^{-if_{N}(t)}\mu_{N}(t)]|{\rm vac}\rangle$ at the time $t$ with the accumulated global phases $f_{1}(t)$ and $f_{N}(t)$, respectively, where $F[\cdot]$ represents an arbitrary function of operators.
In parallel, for the systems governed by the Hermitian-conjugate Hamiltonian $H_{\rm rot}^{\dagger}(t)$ in the dual space, the upper triangularization condition (10) can activate $\mu_{1}(t)$ and $\mu_{N}^{\dagger}(t)$ as nonadiabatic passages in the Heisenberg picture. The Hermitian conjugate of Eq. (11) takes the down triangularization form of
@@
$\mu_{1}(0)\rightarrow e^{if_{1}^{*}(t)}\mu_{1}(t),\quad\mu_{N}^{\dagger}(0)\rightarrow e^{-if_{N}^{*}(t)}\mu_{N}^{\dagger}(t)$ (19)

### v2

@@ Abstract
Current studies about the non-Hermitian continuous-variable systems heavily revolved around the singularities in the eigen-spectrum by mimicking their discrete-variable counterparts. The growing discussions over the nonunitary features in the time evolution, however, are limited in scalability and controllability. We here develop a general theory to control an arbitrary number of bosonic modes under the time-dependent non-Hermitian Hamiltonian. Far beyond the subspace of few excitations, our control theory operates in the full Hilbert space within the Heisenberg framework and exploits the gauge potential underlying the instantaneous frames rather than the eigen-spectrum. In particular, the instantaneous frames are defined by time-dependent ancillary operators as linear combinations of the laboratory-frame operators; while the associated gauge potential arises from the unitary transformation connecting the time-dependent and stationary ancillary frames. We find that the upper triangularization condition of the non-Hermitian Hamiltonian’s coefficient matrix in the stationary ancillary frame gives rise to the nonadiabatic passages of two time-dependent ancillary operators, also leading to the exact solutions of the time-dependent Schrödinger equation. At the end of these passages, the probability conservation of the system wavefunction can be automatically restored without artificial normalization. Our theory is exemplified with the perfect and nonreciprocal state transfers in a cavity magnonic system. The former holds for arbitrary initial states and is irrelevant to both parity-time symmetry of the coefficient matrix and exceptional points of the eigen-spectrum; and the latter is consistent with the coherent perfect absorbtion. We essentially constructs the universal quantum control (UQC) theory for the non-Hermitian continuous-variable systems, promising a powerful and reliable approach for their coherent control.
@@ I Introduction
Conventional control protocols for non-Hermitian systems are spectrum-dependent, surrounding such as the broken or unbroken phases of parity-time ($\mathcal{PT}$) symmetry Bender and Boettcher (1998); Ashida et al. (2020); Bender and Hook (2024) of Hamiltonian or scattering matrix Metelmann and Clerk (2015); Ramezani et al. (2014); Longhi (2015); Jin et al. (2016); Ramezani et al. (2016); Huang et al. (2017); Jin and Song (2018); Harder et al. (2018); Grigoryan et al. (2018); Wang et al. (2019a); Xu and Jin (2024) and the presence or absence of the exceptional points (EP) Heiss (2012). $\mathcal{PT}$-symmetry of Hamiltonian provides a relaxed criterion for determining whether the eigen-spectrum of system is real or complex Bender and Boettcher (1998). EPs describe the coalescence of eigenvalues and eigenstates Heiss (2012), at which the biorthogonal condition breaks down Brody (2013). Spectrum-based protocols often impose stringent experimental requirements Ibáñez et al. (2011); Torosov et al. (2013, 2014). More seriously, the extraction of eigen-spectrum becomes increasingly challenging when the quantum system extends from time-independent to time-dependent and from few dimensions to a large scale. The current work aims to establish a universal theory for non-Hermitian continuous-variable systems that does not mimic their discrete-variable counterparts Bender and Boettcher (1998); Ashida et al. (2020); Bender and Hook (2024); Jin and Jing (2025a). It also constitutes a nontrivial extension of the universal quantum control (UQC) Jin and Jing (2025b, c, d, e, f), beyond the closed systems Jin and Jing (2025g) isolated from the external environment.
@@
Non-Hermitian Hamiltonian is the ubiquitous description for continuous-variable systems, such as photonic systems Guo et al. (2009); Regensburger et al. (2012); Ge and Türeci (2013); Peng et al. (2014); Hodaei et al. (2017); Zhang et al. (2020); Bergman et al. (2021), acoustic systems Zhu et al. (2014); Fleury et al. (2015); Christensen et al. (2016); Liu et al. (2018); Wang et al. (2019b); Tang et al. (2020), and cavity magnonic systems Zhang et al. (2017, 2019); Zhao et al. (2020); Wang et al. (2022); Qian et al. (2024); Yang et al. (2024); Zhang et al. (2025). Typically it arises either from the unavoidable energy or material exchange with the environment Metelmann and Clerk (2015); Ramezani et al. (2014); Longhi (2015); Jin et al. (2016); Ramezani et al. (2016); Huang et al. (2017); Jin and Song (2018); Harder et al. (2018); Grigoryan et al. (2018); Wang et al. (2019a); Xu and Jin (2024), or from post-selection over no-quantum-jump trajectories Han et al. (2024), leading to the nonunitary evolution and the violation of probability conservation of the system wavefunction Ashida et al. (2020). Owing to the cooperative coupling to a common reservoir Metelmann and Clerk (2015); Wang et al. (2019a), the non-Hermitian Hamiltonian accounts for the gain or loss effects associated with individual bosonic modes Zhang et al. (2017, 2025) and the dissipative coupling between the system modes Metelmann and Clerk (2015); Harder et al. (2018); Grigoryan et al. (2018); Wang et al. (2019a). In practice, the loss effect on a cavity mode can be effectively converted to the gain effect by applying external microwave fields to its ports Zhang et al. (2017, 2025).
In analog to the two-level systems Ibáñez et al. (2011); Torosov et al. (2013, 2014); Chen et al. (2016); Luan et al. (2022), the shortcut-to-adiabaticity technique was developed to describe the non-Hermitian cavity magnonic systems Zhang et al. (2022) in a highly restrictive subspace. Nevertheless, it is not an approach for the full Hilbert space. Scattering matrix based on the input-output theory Zhang et al. (2017); Ramezani et al. (2014); Longhi (2015); Metelmann and Clerk (2015); Jin et al. (2016); Ramezani et al. (2016); Huang et al. (2017); Jin and Song (2018); Harder et al. (2018); Grigoryan et al. (2018); Wang et al. (2019a); Xu and Jin (2024) formulates a phenomenological and yet rough description for the non-Hermitian continuous-variable systems. Through precise match between coherent and engineered dissipative interaction, this approach can be used to demonstrate the directional quantum amplification Metelmann and Clerk (2015), the unidirectional microwave invisibility Jin and Song (2018); Wang et al. (2019a), and coherent perfect absorption Ramezani et al. (2014); Longhi (2015); Jin et al. (2016); Ramezani et al. (2016); Huang et al. (2017); Jin and Song (2018); Xu and Jin (2024). Most existing works are constrained by the system size and the controllability of target states, excessively depend on the system spectrum, have to repeatedly renormalize the wavefunction during the time evolution, and then cannot be regarded as a systematic and versatile theoretical framework for engineering the non-Hermitian bosonic systems.
@@
In contrast to the eigen-spectrum-based approaches, the universal quantum control framework Jin and Jing (2025b, c, d, e, a, f, g) can provide a broader and more general perspective by exploiting the gauge potential underlying the differential manifold of rotating frames. With the time-dependent Hamiltonian under certain conditions, a UQC theory has been formulated for non-Hermitian discrete-variable systems Jin and Jing (2025a), activating a single nonadiabatic passage in either bra or ket space. In this paper, it is extended to the multiple-mode bosonic systems governed by the non-Hermitian and time-dependent Hamiltonian that is obtained by the Lindblad master equation with the quantum-jump terms retained. Transformation from the time-dependent to the time-independent (stationary) ancillary frames introduces a gauge potential, by which the triangularization condition of the Hamiltonian’s coefficient matrix enables the exact solutions of the Heisenberg equation for two of the ancillary operators. They active two system passages in both bra and ket spaces toward arbitrary target modes. In contrast to the conventional treatments of non-Hermitian systems Daley et al. (2009); Uzdin et al. (2012); Ibáñez and Muga (2014); Ashida and Ueda (2018); Dóra and Moca (2020), our theory is objective, deterministic, and scalable in terms of control target and inherently conserves probability at the end of the passage without artificial normalization. In a cavity magnonic system with gain or loss effects on individual modes and a dissipative coupling between them, it is found that arbitrary state in the cavity mode can be perfectly transferred to the magnon mode and unidirectional perfect absorbtion Ramezani et al. (2014); Longhi (2015); Jin et al. (2016); Ramezani et al. (2016); Huang et al. (2017); Jin and Song (2018); Xu and Jin (2024) can be realized at the operational level if these two modes are exchanged.
@@
The rest of this paper is structured as follows. In Sec. II, we provide a general theory for solving the Heisenberg equation for the bosonic operators driven by a non-Hermitian and time-dependent Hamiltonian. The Hamiltonian’s coefficient matrix can be triangularized in a stationary frame and then yields two ancillary operators to unravel Heisenberg passages. In Sec. III.1, the general theory is applied to the non-Hermitian cavity-magnonic system; and in Sec. III.2, we construct the relevant Heisenberg passages that in general do not conserve the state-norm. Section IV.1 demonstrates that an arbitrary state prepared in the cavity mode can be perfectly transferred to the magnon mode through a designed passage, irrelevant to the parity-time-symmetry and the presence or absence of exceptional points. In addition, Sec. IV.2 shows that under the same setting a unidirectional perfect absorbtion occurs when the two modes are exchanged. The entire work is summarized in Sec. V. Appendix A provides a brief recipe for constructing the ancillary operators as well as the unitary transformation between the time-dependent and stationary ancillary frames. Appendix B details the derivation of the non-Hermitian Hamiltonian based on the Lindblad master equation for the cavity-magnonic system.
@@ II General framework
where $\vec{a}\equiv(a_{1},a_{2},\cdots,a_{N})$ and $\vec{a}^{\dagger}=(a_{1}^{\dagger},a_{2}^{\dagger},\cdots,a_{N}^{\dagger})$ are row vectors of operator, and $H^{a}(t)$ is an $N\times N$ time-dependent coefficient matrix with Hermitian-conjugate $H^{a}(t)]^{\dagger}$. $H^{a}(t)\neq[H^{a}(t)]^{\dagger}$. The superscript $T$ denotes the matrix transposition, also transforming a row vector to its corresponding column vector. Due to the noncommunity, non-Hermicity, and the infinite-dimensional Hilbert space of the time-dependent Hamiltonian, solving the Schrödinger equation for non-Hermitian continuous-variable systems is generally challenging. However, our UQC theory Jin and Jing (2025b, c, d, e, a, f, g) can be extended to the non-Hermitian bosonic systems, providing a fundamental framework to partially solve Eqs. (1a) and (1b) in the same time.
Without loss of generality, we start with the system dynamics in the ket space, as given by Eq. (1a) with $H(t)$ in Eq. (2). In the spirit of universal quantum control for Hermitian bosonic systems Jin and Jing (2025g), the system dynamics can be described in the ancillary representation associated with a completed set of time-dependent ancillary operators $\mu_{k}(t)$’s, $1\leq k\leq N$. $\{\mu_{k}(t)\}$ are superposed by the laboratory-frame bosonic operators $\{a_{k}\}$ with an $N\times N$ unitary transformation matrix $\mathcal{M}^{\dagger}(t)$:
@@
A specific unitary transformation matrix $\mathcal{M}^{\dagger}(t)$ and the relevant ancillary operators are constructed in appendix A. Due to Eq. (3) and the unitary feature of $\mathcal{M}^{\dagger}(t)$, one can confirm that the ancillary operators $\mu_{k}(t)$’s satisfy the canonical communication relation, i.e., $\mu_{j}(t),\mu_{k}^{\dagger}(t)]=\delta_{jk}$. Using Eq. (3), the time-dependent Hamiltonian (2) can be formulated as
@@
is the Hamiltonian’s coefficient matrix in the representation of time-dependent ancillary bases $\mu_{k}(t)|{\rm vac}\rangle$ with $|{\rm vac}\rangle$ the vacuum state.
To proceed, we consider the rotation from the time-dependent ancillary operators to the time-independent or stationary ancillary operators, i.e., $\vec{\mu}_{t}\rightarrow\vec{\mu}_{0}$ with $\vec{\mu}_{0}=[\mu_{1}(0),\mu_{2}(0),\cdots,\mu_{N}(0)]$. Such a rotation can be achieved by the unitary transformation $\mathcal{V}(t)$ as $\mathcal{V}^{\dagger}(t)\mu_{k}(t)\mathcal{V}(t)\rightarrow\mu_{k}(0)$, where $\mathcal{V}(t)$ is determined by $\mathcal{M}^{\dagger}(t)$ in Eq. (3) (see appendix A for details). Generally, in the rotating frame with respect to $\mathcal{V}(t)$, we have
@@
Consequently, the time-dependent Schrödinger equations (1) are transformed as
@@
The preceding rotation to the stationary representation (6) does not directly relieve the difficulty in solving the non-Hermitian Schrödinger equation (7). On one hand, it is not appropriate to apply the commutation condition about the Hamiltonian’s coefficient matrix and the projection operator in the representation of time-independent ancillary modes established for the Hermitian bosonic network Jin and Jing (2025g), since the non-Hermitian coefficient matrix $H^{\mu}(t)-\mathcal{A}(t)$ is generally non-diagonalizable, i.e., $H^{\mu}_{km}(t)\neq H^{\mu}_{mk}(t)$ for $k\neq m$. On the other hand, the triangularization condition about the rotated Hamiltonian for discrete-variable systems Jin and Jing (2025a) also fails here due to the fundamentally distinct statistics of bosonic modes.
Main result.— we prove that the upper triangularization of the coefficient matrix of $H_{\rm rot}(t)$ is a sufficient condition for activating two useful Heisenberg-picture passages, which is only relevant to the differential manifold rather than the spectral characteristics. Specifically, the upper triangularization condition for the coefficient matrix $H^{\mu}(t)-\mathcal{A}(t)$ is given by Jin and Jing (2025a); Ju et al. (2024)
@@
where $\Pi^{k}$, $1\leq k\leq N$, is the projection operator or an $N\times N$ matrix defined by $\Pi^{k}_{jm}=\delta_{jk}\delta_{mk}$.
@@
The dynamics of an arbitrary operator $\tilde{\mathcal{O}}_{S}$ can be obtained by $\mathcal{O}_{H}(t)=V_{\rm rot}^{\dagger}(t)\tilde{\mathcal{O}}_{S}U_{\rm rot}(t)$ according to the non-Hermitian Heisenberg equation Miao and Xu (2016) with the evolution operators given by Eq. (9). Under the Hamiltonian (11), the dynamics of the ancillary operators $\mu_{1}^{\dagger}(0)$ and $\mu_{N}(0)$ are found to be decoupled from the others and therefore can be derived as $v_{1}^{\dagger}(t)\equiv V_{\rm rot}^{\dagger}(t)\mu_{1}^{\dagger}(0)U_{\rm rot}(t)$ and $v_{N}(t)\equiv V_{\rm rot}^{\dagger}(t)\mu_{N}(0)U_{\rm rot}(t)$. In particular, we have
@@
Equations (16) and (17) indicate that the ancillary operators $\mu_{1}^{\dagger}(t)$ and $\mu_{N}(t)$ are activated as useful passages toward the desired target mode. The systems initially prepared in the states $F[\mu_{1}^{\dagger}(0)]|{\rm vac}\rangle$ and $F[\mu_{N}(0)]|{\rm vac}\rangle$ will evolve to $F[e^{if_{1}(t)}\mu_{1}^{\dagger}(t)]|{\rm vac}\rangle$ and $F[e^{-if_{N}(t)}\mu_{N}(t)]|{\rm vac}\rangle$ at the time $t$ with the accumulated global phases $f_{1}(t)$ and $f_{N}(t)$, respectively, where $F[\cdot]$ represents an arbitrary function of operators.
In parallel, for the systems governed by the Hermitian-conjugate Hamiltonian $H_{\rm rot}^{\dagger}(t)$ in the dual space, the upper triangularization condition (10) can activate $\mu_{1}(t)$ and $\mu_{N}^{\dagger}(t)$ as nonadiabatic passages in the Heisenberg picture. The Hermitian conjugate of Eq. (11) takes the lower triangularization form of
@@
$\mu_{1}(0)\rightarrow e^{-if_{1}^{*}(t)}\mu_{1}(t),\quad\mu_{N}^{\dagger}(0)\rightarrow e^{if_{N}^{*}(t)}\mu_{N}^{\dagger}(t)$ (19)